OpenStudy (anonymous):

Which value of x solves the equation cos x° = sin (20° + x°), where 0 < x < 90?

OpenStudy (anonymous):

A: 30° B: 35° C: 40° D: 45° E: 55° @iGreen. @AnswerMyQuestions @Legends

OpenStudy (anonymous):

I found the solution online but I'm not quite sure how they got it; sin(90 - x) = sin(20 + x) 90 - x = 20 + x 2x = 70 x = 35 degs